Inhaltsverzeichnis

William Wordsworth, The Daffodils . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 2

William Wordsworth, The Daffodils  2/2

William Wordsworth,

The Daffodils

I wander’d lonely as a cloud That floats on high o’er vales and hills, When all at once I saw a crowd, A host of golden daffodils, Beside the lake, beneath the trees Fluttering and dancing in the breeze.

Continuous as the stars that shine And twinkle on the milky way, They stretch’d in never-ending line Along the margin of a bay: Ten thousand saw I at a glance Tossing their heads in sprightly dance.

The waves beside them danced, but they Out-did the sparkling waves in glee: - A poet could not but be gay In such a jocund company: I gazed - and gazed - but little thought What wealth the show to me had brought.

For oft, when on my couch I lie In vacant or in pensive mood, They flash upon that inward eye Which is the bliss of solitude; And then my heart with pleasure fills And dances with the daffodils.

Verfasser dieses englischen Gedichtes ist William Wordsworth (∗1770 – †1850).

Arg-39-73 (2003-10-17 22:02:21)

Origina anglalingva teksto de WILLIAM WORDSWORTH (1770 - 1850)

http://www.poezio.net/version?poem-id=39&version-id=73